OpenStudy (anonymous):

6(n+5)=66 n=6 how does it equal 6?

OpenStudy (anonymous):

multiply in the 6 6n+30=66 then u can take it from there

OpenStudy (anonymous):

so divide both sides by 6. and subtract 5 to both sides. imagine unwrapping a gift.

OpenStudy (anonymous):

Distribute. 6n + 30 = 66 6n = 66 - 30 6n = 36 n = 6.

OpenStudy (anonymous):

6(6+5)=66 6*11=66 66=66